OpenStudy (anonymous):

Solve the equation and check for extraneous solutions. Click to see question!

OpenStudy (anonymous):

\[\left( -2-x \right)^{0.5}-x=2\]

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

(-2-x)^0.5-x = 2 sqrt(-2-x) = x +2 -2-x = x^2+4x+4 x^2+5x+6 = 0 Factor: (x+3)(x+2) = 0 x = -2 or x = -3 (-2-x)^0.5-x = 2 Check x = -2: (-2+2)^(0.5) +2 = 2 sqrt(0) + 2 = 2 2 = 2 Check x = -3 (-2+3)^(0.5) + 3 = 2 sqrt(1)+3 = 2 4 = 2 False --- Only solution: x = -2

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Yeah I'm getting there! Appreciate your help!

OpenStudy (anonymous):

math is very simple ;) only logic. :)

OpenStudy (anonymous):

Yeah but can you answer my other question cause your the only that explain it good and nobody else can answer it.

OpenStudy (anonymous):

Pretty pleaseee!